Liz and Bob just had a baby named Isabelle, and they want to save enough money for Isabelle to go to college. Assume that they start making monthly payments when Isabelle is 33 into an ordinary annuity earning 3.75%3.75%, and they calculate that they will need $27,900.00$27,900.00 by the time Isabelle turns 1818. How much should they deposit every month so that they reach their goal?
Answer #1

FV = PMT \frac{(1 + i)^n - 1}{i}

i = 0.0375/12

n = 12*(18-3) = 180

27900 = PMT * ((1 + 0.0375/12)^180 - 1)/(0.0375/12)

27900 = PMT *241.12512

hence

PMT = 27900/241.12512= 115.7075